OpenStudy (anonymous):

1. [2.01] a = a is an example of which algebraic property? Symmetric Property Transitive Property Reflexive Property Commutative Property of Multiplication 2. [2.01] 3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property 3. [2.01] Solve: x − 2 = −6 −8 8 −4 4 4. [2.01] Solve: = 10 5 2 50 5. [2.02] Solve: + 5 = 12 −14 14 34 −34 6. [2.02] Solve: −2(x − 8) = −10 −13 13 −3 3 7. [2.02] Solve: 8x − 10 + 18 = 24 4 −2 6.5 2 8.

OpenStudy (anonymous):

1. [2.01] a = a is an example of which algebraic property? Symmetric Property Transitive Property Reflexive Property Commutative Property of Multiplication 2. [2.01] 3(a + b) = 3a + 3b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property 3. [2.01] Solve: x − 2 = −6 −8 8 −4 4 4. [2.01] Solve: = 10 5 2 50 5. [2.02] Solve: + 5 = 12 −14 14 34 −34 6. [2.02] Solve: −2(x − 8) = −10 −13 13 −3 3 7. [2.02] Solve: 8x − 10 + 18 = 24 4 −2 6.5 2 8. [2.02] Solve: 5(x − 1) − (x + 5) = 2 3 2 1 9. [2.03] Solve: 4(2x − 1) − 3 = 9(x − 4) −29 32 13 29 10. [2.03] Solve: 2(x + 1) = 2x + 5 Any Real Number No Solution zero 3 11. [2.03] Solve: 8(x − 1) = 6(x + 2) − 4 zero 8 3 12. [2.05] The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. Age Ticket Price Under three years free Three to twelve years $32 Over twelve years $44 On a certain day, the company took 82 people on whale watching trips. There were 54 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (54 + x)32 + (82 − (54 + x)) ∙ 44 = C (54 − x)32 + (82 − 54) ∙ 44 = C (82 + x)32 + (82 − (54 + x)) ∙ 44 = C (82 − x)32 + (82 − 54) ∙ 44 = C 13. [2.05] The sum of two consecutive even integers is 226. What is the larger integer? 114 116 112 118 14. [2.05] In a certain county, the number of Charter Schools is ten less than twice the number of Alternative Schools. We know that there are 46 Charter Schools in the county. How many Alternative Schools are in the county? 28 56 18 40 15. [2.05] Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter? 10 18 12 20 16. [2.05] Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now? 20 10 30 15 17. [2.06] Solve for x: = 6. 13 14 23 10 18. [2.06] Solve for x: x + 2 = 5. 28 20 14 12 19. [2.06] Solve for x: (x − 4) = 2x. −8 1 −1 20. [2.08] Solve for L: A = L • W • H L = L = L = L = 21. [2.08] The formula for degrees Celsius is: C = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

1) reflexive property 2) Distributive Property 3) x=-4